OpenStudy (anonymous):

each function is either even or odd .use f(-x) to state which situation applies f(x)=5x^3-8x please help

OpenStudy (anonymous):

do you get the difference between even and odd functions?

OpenStudy (anonymous):

This is a good resource to start with: http://www.purplemath.com/modules/fcnnot3.htm Feel free to ask if you still have any questions

OpenStudy (unklerhaukus):

Find f(-x) :

OpenStudy (anonymous):

i will look at it thank you

OpenStudy (unklerhaukus):

\[f(x)=5x^3-8x\] so \[f(-x)=5(-x)^3-8(-x)\] .. can you simplify this?

OpenStudy (anonymous):

-5x^3+8x

OpenStudy (unklerhaukus):

good, so you found \[f(-x)=-5x^3+8x\] if we now factor out the negative sign\[f(-x)=-(5x^3-8x)\] We have the negative of the original function\[f(-x)=-f(x)\] This is the definition of an odd function.

OpenStudy (anonymous):

thank you :)

OpenStudy (unklerhaukus):

if we had gotten\[f(-x) = f(x)\], (after simplification), it would have been an even function